OpenStudy (anonymous):

Rewrite with only sin x and cos x. sin 2x - cos 2x

OpenStudy (anonymous):

- 2 sin^2(x)+sin 2x - 1

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

cos 2x=cos^2(x) - sin^2(x) = 2 cos^2(x) - 1 = 1 - 2 sin^2(x)

OpenStudy (anonymous):

A. 2 sinx cosx - 1 + 2 sin2x B. 2 sin x cos2x - 1 + 2 sin2x C. 2 sin x cos2x - sin x + 1 - 2 sin2x D. 2 sin x cos2x - 1 - 2 sin2x These are my chocies and I don't know If your wrong or not

OpenStudy (anonymous):

sin2x=2sinx.cosx cos2x=cos^2x-sin^2x sin2x-cos2x= 2sinx.cosx-cos^2x+sin^2x

OpenStudy (anonymous):

a

OpenStudy (anonymous):

right

OpenStudy (anonymous):

cos^2x=1-sin^2x convert it

OpenStudy (anonymous):

Thank you